# Thermodynamic Cost of Random-Time Protocols

### v3

@@ I Introduction
In this paper, we identify a thermodynamic cost associated with randomly timed protocols that is informational in nature. Taking this additional cost into consideration in setups that use random times, we find that the second law of thermodynamics is upheld, and hence we obtain a thermodynamically consistent setup.
Specifically, drawing an analogy with Maxwell’s demon in Szilard’s engine [27, 28], we argue that a non-autonomous demon implementing a random-time protocol must store the time variable in a memory device. The associated thermodynamic cost arises, as with Szilard’s engine, from restoring the memory device to its initial state [27, 28]. Indeed, measurements increase the entropy of the device on which the outcome of the measurement is stored. Erasing the information stored on the memory device is a reduction of entropy that requires useful work [29]. As we demonstrate, apparent violations of the second law of thermodynamics in randomly timed protocols – such as in stochastically stopped processes [1, 2, 3, 4, 5, 6], isothermal cyclic protocols  [7, 8, 9, 10, 11] or heat engines surpassing the Carnot efficiencies [12] – are resolved once we include in these setups a memory device. Since the need for a memory arises from the random nature of the timing of the external protocol, our results apply generically to external protocols relying on random times. For example, stochastic resetting protocols do not only require input work to move the particle, as calculated in Refs. [25, 26, 25], but also require input work to erase the memory device that stores the random time.
@@
Furthermore we introduce in this paper a non-autonomous demon that directly measures first-passage times. Although first-passage demons have been considered before in the literature, see Refs. [7, 8, 9, 10, 11], these non-autonomous demons in fact measure the whole trajectory of the system up to the first-passage time through continuous monitoring. In setups of non-autonomous demons, continuous monitoring requires a large number of measurements, as continuous monitoring is implemented through a protocol that measures the state of the system at regularly spaced time intervals. Here, instead, we develop a simpler setup for which a non-autonomous demon directly measures the first-passage time using a single measurement. We present concrete examples of setups that measure first-passage times directly without continuous monitoring.
@@
The paper is structured as follows. In Sec. II, we introduce the system setup for externally controlled protocols that operate with random times, focusing on the isothermal case. In Sec. III, we argue, by analogy with Szilard’s engine, that a non-autonomous demon implementing a random-time protocol requires a memory device to run the protocol. In Sec. IV, we show that the storage of a random time is sufficient to derive the second law of thermodynamics for systems controlled by random-time protocols, further confirming that a memory device is needed in these setups. As we demonstrate, the second law of thermodynamics follows from combining Landauer’s erasure principle [42, 27, 28] with the formalism of information thermodynamics for non-autonomous Maxwell demons with a path-dependent protocol [38, 39, 40]. The Secs. V to VII explore a paradigmatic example of a temporal information engine, which is the random-time version of Szilard’s engine as illustrated in Fig. 1. This model is arguably the simplest example of a temporal information engine. It is a simplification of previous setups based on continuous Maxwell demons [7, 8, 9, 10, 11], as the demon measures the random time directly without continuous monitoring. In Sec. V, we consider the random-time Szilard engine in discrete time and without measurement errors, and in Sec. VI we consider the corresponding model in continuous time. In Sec. VII we develop a simple model for random-time detection without continuously monitoring the system, confirming that the model is physically sound. In Sec. VIII we show that stochastic heat engines that use random-time protocols satisfy the Carnot bound on their efficiency, which is further evidence in support of the claim that memory storage is necessary in random-time protocols. Section IX compares the setups of Maxwell demons that continuously monitor a system with those that directly measure first-passage times. We end the paper with a discussion in Sec. X and a few appendices.
@@ IV Second law of thermodynamics for non-autonomous demons that use random times
is the Shannon entropy of the stopping time $T$ measured in nats; here $p_{T}$ is the probability distribution of $T$. The inequality (7) states that random-time protocols can thermodynamically perform work on an external load, i.e., $\langle W\rangle<0$, but for cyclic protocols with $\Delta f=0$ the work is bounded by
@@
We note that the Eq. (7) is a stochastic thermodynamics result. It can be derived in the setup of Markov processes that satisfy local detailed balance (see Appendix A). An alternative derivation of the inequality (7) that is based on continuously monitoring the system is presented in Ref. [10]. This is because at the level of mathematical formulation, both the sleeping demon and the conventional continuous demon can be described in a unified setup of error-free non-autonomous Maxwell demon with a path-dependent protocol $\lambda(t;X^{t}_{0})$ [38, 39, 40]. Furthermore, following Ref. [41], the inequality (7) can be further refined by taking into account the unavailable information. However, our purpose is to derive the second law of thermodynamics for randomly timed protocols, and for this purpose the inequality (7) provides the most parsimoniuous explanation.
@@ V Random-time Szilard engine: discrete version
Figure 2: Random-time Szilard engine in discrete time without measurement error; (see Sec. V). Comparison between the work $\beta\langle W\rangle=\ln(\ell)$ (blue circles) done by the Szilard engine on the external load and the entropy $H(T)$ (green squares) of the stored time $T$ as a function of the length $\ell$ of the lattice; the hopping probability $p_{0}=1$. The (orange) star denotes the value $H(T)=2\ln 2$ for $\ell=2$ (see Appendix B), and the red dotted line plots the function $\ln\ell^{2}$.
@@ VI Random-time Szilard engine: continuous version
Figure 3: Random-time Szilard engine in continuous time without measurement error; (see Sec. VI). The ratio between the work $\beta\langle W\rangle$ done by the system on the external load and the entropy $H(T)$ as a function of the length ratio $\tilde{\ell}$$\ell$ for three given values of $\zeta=D\theta/\ell^{2}$. See Appendix C for explicit expressions of the plotted functions.
@@ VII Random-time detectors
Figure 4: Random-time Szilard engine with measurement error; (see Sec. VII). Average work that a random-time Szilard engine exerts on an external load as a function of the parameter $\pi_{\ell}/p_{0}$ that characterises the quality of the detector. The other detector parameters are $\pi_{0}=0.0001$ and $\beta e_{\uparrow}=4$. The hopping probability $p_{0}=0.1$ and $\ell$ is as given in the legend. The average work is obtained from the formula $\beta\langle W\rangle=(2p_{+}-1)\ln(\ell-1)$, where $p_{+}$ is the probability that $X(T)=\ell$; this is Eq. (15) with $\tilde{\ell}=1$. Each marker is an average over $1e+6$ simulated samples.
@@ VIII.2 Analysis with memory storage: recovering the Carnot bound
Using (29) in (28) we recover Carnot’s bound (18). In conclusion, a cyclic heat engine satisfies the Carnot bound, even if it uses a protocol with a random time $T$.
IX Measuring first-passage times: comparison between direct methods and continuous monitoring
@@ X Discussion
Consequently, using random times in external protocols incurs a thermodynamic cost that is informational in nature. This additional cost must also be accounted for in protocols not typically associated with Maxwell’s demons. A pertinent example is stochastic resetting, a process that maintains a nonequilibrium steady state by returning a diffusing particle to the origin at a random time [13, 14]. This process requires an input of work, as energy must be expended to bring the particle back to the origin [25]. However, there is an additional cost related to storing the random time $T$ in memory. Neglecting this cost can lead to apparent violations of the second law of thermodynamics.
To illustrate the consequences of temporal information on nonequilibrium thermodynamics, we have introduced a non-autonomous Maxwell demon that directly measures the random time $T$, and which we called the sleeping demon. We have developed setups that measure first-passage times directly, unlike the conventional approach based on continuous monitoring that is typically employed in the literature [10, 7, 8, 9, 11]. In the direct approach, the cost of measurement is fixed and independent on the size of the system. This is different from setups based on continuous measurements, for which multiple measurements are performed, and therefore the cost of measurement grows with the system size.
@@ Acknowledgements.
Appendix B Entropy $H(T)$ for the random-time Szilard engine in discrete time
@@ C.2 Probability distribution of the first-passage time $T_{\rm fp}$
Notice that $p_{T_{\rm fp}}$ is normalised (see Appendix D), and that $T_{\rm fp}$ scales as $\ell^{2}/D$.
@@ C.3 The entropy $H(T)$ of the random control time $T$
Figure 6: Entropy $H(T)$ of the distribution $p_{T}(t)$, given by Eq. (54), as a function of $\zeta$. The dashed line denotes $\ln(\zeta)$ for comparison.

### v5

@@ I Introduction
In this paper, we identify a thermodynamic cost associated with randomly timed protocols that is informational in nature. Specifically, drawing an analogy with Maxwell’s demon in Szilard’s engine [27, 28], we argue that a non-autonomous demon implementing a random-time protocol must store the time variable in a memory device. The associated thermodynamic cost arises, as with Szilard’s engine, from restoring the memory device to its initial state [27, 28]. As we show, apparent violations of the second law of thermodynamics are resolved once this additional cost is taken into account.
@@
Furthermore we introduce in this paper a non-autonomous demon that directly measures first-passage times. Although first-passage demons have been considered before in the literature, see Refs. [7, 8, 9, 10, 11], these non-autonomous demons measure the whole trajectory of the system up to the first-passage time through continuous monitoring. For non-autonomous demons, this approach requires a large number of measurements, as the state of the system should be measured at regularly spaced intervals of time. Here, instead, we develop a simpler setup for which a non-autonomous demon directly measures the first-passage time using a single measurement. We present concrete examples of setups that directly measure first-passage times.
@@
The paper is structured as follows. In Sec. II, we introduce the system setup for externally controlled protocols that operate with random times, focusing on the isothermal case. In Sec. III, we argue, by analogy with Szilard’s engine, that a non-autonomous demon implementing a random-time protocol requires a memory device to run the protocol. In Sec. IV, we show that storage of the random time is sufficient to derive the second law of thermodynamics for isothermal systems controlled by random-time protocols, further confirming that a memory device is needed in these setups. The derivation uses Landauer’s erasure principle [42, 27, 28] and information thermodynamics for non-autonomous Maxwell demons with a path-dependent protocol [38, 39, 40]. The Secs. V to VII explore a paradigmatic example of a temporal information engine, which is a random-time version of Szilard’s engine, as illustrated in Fig. 1. This model is arguably the simplest example of a temporal information engine. It is a simplification of previous setups based on continuous Maxwell demons [7, 8, 9, 10, 11], as the demon measures the random time directly without continuous monitoring. In Sec. V, we consider the random-time Szilard engine in discrete time and without measurement errors, and in Sec. VI we consider the corresponding model in continuous time. In Sec. VII we develop simple models for random-time detection without continuously monitoring the system, confirming that first-passage time measurements are physically sound. In Sec. VIII we show that stochastic heat engines that use random-time protocols satisfy the Carnot bound if the random time is stored in a memory. Section IX compares the setups of Maxwell demons that continuously monitor a system with those that directly measure first-passage times. We end the paper with a discussion in Sec. X and a few appendices that contain the derivations of results presented in the main text.
@@ IV Second law of thermodynamics for non-autonomous demons that use random times
is the Shannon entropy of the stopping time $T$ measured in nats; here $p_{T}$ is the probability distribution of $T$. The inequality (7) states that random-time protocols can thermodynamically perform work on an external load, i.e., $\langle W\rangle<0$. For cyclic protocols with $\Delta f=0$ the work is bounded by
@@
We note that the Eq. (7) is a stochastic thermodynamics result. It can be derived in the setup of Markov processes that satisfy local detailed balance (see Appendix A). An alternative derivation of the inequality (7) that is based on continuously monitoring the system is presented in Ref. [10]. This is because at the level of mathematical formulation, both the sleeping demon and the conventional continuous demon can be described in the unified setup of error-free non-autonomous Maxwell demon with a path-dependent protocol $\lambda(t;X^{t}_{0})$ [38, 39, 40]. Furthermore, following Ref. [41], the inequality (7) can be further refined by taking into account the unavailable information. However, our purpose is to derive the second law of thermodynamics for randomly timed protocols, and for this purpose the inequality (7) provides the most parsimoniuous explanation.
@@ V Random-time Szilard engine: discrete version
Figure 2: Random-time Szilard engine in discrete time without measurement error (see Sec. V). Comparison between the work $\beta\langle W\rangle=\ln(\ell)$ (green squares) done by the Szilard engine on the external load and the entropy $H(T)$ (blue circles) of the stored time $T$ as a function of the length $\ell$ of the lattice; the hopping probability $p_{0}=1$. The blue circles denote the entropy $H(T)$ of the probability mass function of $T={\rm min}\left\{t\geq 0:X(t)=\ell\right\}$ obtained from repeated simulations of the random walk process with a uniform initial distribution for $X(0)$. The (orange) star denotes the value $H(T)=2\ln 2$ for $\ell=2$ (see Appendix B), and the red dotted line plots the function $\ln\ell^{2}$.
@@ VI Random-time Szilard engine: continuous version
Figure 3: Random-time Szilard engine in continuous time without measurement error (see Sec. VI). The ratio between the work $\beta\langle W\rangle$ done by the system on the external load and the entropy $H(T)$ as a function of the length ratio $\tilde{\ell}$$\ell$ for three given values of $\zeta=D\theta/\ell^{2}$. See Appendix C for explicit expressions of the plotted functions.
@@ VII Random-time detectors
Figure 4: Random-time Szilard engine with measurement error (see Sec. VII). Average work that a random-time Szilard engine exerts on an external load as a function of the parameter $\pi_{\ell}/p_{0}$ that characterises the quality of the detector. The other detector parameters are $\pi_{0}=0.0001$ and $\beta e_{\uparrow}=4$. The hopping probability $p_{0}=0.1$ and $\ell$ is as given in the legend. The average work is obtained from the formula $\beta\langle W\rangle=(2p_{+}-1)\ln(\ell-1)$, where $p_{+}$ is the probability that $X(T)=\ell$; this is Eq. (15) with $\tilde{\ell}=1$. Each marker is an average over $1e+6$ simulated samples.
@@ VIII.2 Analysis with memory storage: recovering the Carnot bound
Using (29) in (28) and the inequality $\langle Q_{\rm h}\rangle<0$, we recover Carnot’s bound (18). In conclusion, a cyclic heat engine satisfies the Carnot bound, even if it uses a protocol with a random time $T$.
IX Measuring first-passage times: comparison between continuous monitoring and direct methods
@@ X Discussion
Consequently, using random times in external protocols incurs a thermodynamic cost that is informational in nature. This additional cost must also be accounted for in protocols not typically associated with Maxwell demons. A pertinent example is stochastic resetting, a process that maintains a nonequilibrium steady state by returning a diffusing particle to the origin at a random time [13, 14]. This process requires an input of work, as energy must be expended to bring the particle back to the origin [25]. However, there is an additional cost related to storing the random time $T$ in memory. Neglecting this cost can lead to apparent violations of the second law of thermodynamics.
To illustrate the consequences of temporal information on nonequilibrium thermodynamics, we have introduced a non-autonomous Maxwell demon that directly measures the random time $T$, and which we called the sleeping demon. We have developed setups that measure first-passage times directly, unlike the conventional approach based on continuous monitoring that is typically used in the literature [10, 7, 8, 9, 11]. In the direct approach, the cost of measurement is fixed and independent on the size of the system, while for continuous measurement the cost of measurement grows with the system size.
@@ Acknowledgements.
Appendix B Entropy $H(T)$ for the random-time Szilard engine in discrete time and for $\ell=2$
@@ C.2 Probability distribution of the first-passage time $T_{\rm fp}$
Notice that $p_{T_{\rm fp}}$ is normalised and that $T_{\rm fp}$ scales as $\ell^{2}/D$ (see Appendix D).
@@ C.3 The entropy $H(T)$ of the random control time $T$
Figure 6: Entropy $H(T)$ of the distribution $p_{T}(t)$ given by Eq. (54) as a function of $\zeta$. The dashed line denotes $\ln(\zeta)$ for comparison.